OpenStudy (anonymous):

A bicycle wheel rotates 500 times in a minute as long as the rider is pedaling. If the rider stops pedaling, the wheel starts to slow down. Each minute it will rotate only 2/3 as many times as in the preceding minute. How many times will the wheel rotate in the 5th minute after the rider's feet leave the pedals? Round your answer to the nearest unit.

hero (hero):

If the rider stops pedaling and we count from there, then in the first minute it will rotate 1st Minute 500× frac{2}{3} 2nd Minute 500×frac{2}{3}×frac{2}{3} 3rd Minute \[500×\frac{2}{3}×\frac{2}{3}×\frac{2}{3}\] etc...

hero (hero):

Yeah, I think you get it. But we still need the amount of times it rotates in the 5th minute.

hero (hero):

So how many times will it rotate in the fifth minute?

OpenStudy (anonymous):

66!

OpenStudy (anonymous):

thanks Hero :)

hero (hero):

Yup!

hero (hero):

What took you so long?

OpenStudy (anonymous):

I had to feed my dog lol I'm sorry